1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

Le loup-garou The Werewolf The Banshee (An Approach)

tradukita de R. Platteau  tradukita de Alexander Gross tradukita de Max Knight

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître

d’école Et lui dit, “ sur moi dis tout, c’est

ton rôle ! ”

A Werewolf, troubled by his

name, Left wife and brood one night

and came To a hidden graveyard to enlist The aid of a long-dead philolo-

gist.

One night, a banshee slunk

away from mate and child, and

in the gloom went to a village teacher’s

tomb, requesting him: “Inflect

me, pray.”

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître

d’école Et lui dit, “ sur moi dis tout, c’est

ton rôle ! ”

“Oh sage, wake up, please don’t

berate me,” He howled sadly, “Just conju-

gate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then was ready.

The village teacher climbed up straight upon his grave stone with

its plate and to the apparition said who meekly knelt before

the dead:

“ - On peut à ton sujet puisque

nous y sommes Se demander encor bien des choses

en somme. Loup-garoù ? mais avant Il faut se dire : loup-garquand ?

“Well, ‘Werewolf’ is your plural

past, While ‘Waswolf’ is singularly

cast: There’s ‘Amwolf’ too, the pre-

sent tense, And ‘Iswolf,’ ‘Arewolf’ in this

same sense.”

“The banSHEE, in the sub-

ject’s place; the banHERS, the posses-

sive case. The banHER, next, is what they call

objective case–and that is all.”

Loup-garcomment ? et puis Bien évidemment loup-garqui ? ” Toutes ces questions semblèrent au

loup bien absconses, Car en fait ce qu’il voulait c’était

les réponses.

“I know that–I’m no mental

cripple– The future form and participle Are what I crave,” the beast re-

plied. The scholar paused–again he

tried:

The banshee marveled at

the cases and writhed with pleasure,

making faces, but said: “You did not add,

so far, the plural to the singular!”

…  …  …

2/2

“ - Point de réponses, dit le puits

de savoir, Etre imaginaire, nul ne peut te voir

! ” A ces mots l’animal éclata tout en

larmes. “ - Oh!, même les baisers de ma

douce moitié “A ‘Will-be-wolf?’ It’s just too

long: ‘Shall-be-wolf?’ ‘Has-been-

wolf?’ Utterly wrong! Such words are wounds beyond

all suture– I’m sorry, but you have no fu-

ture.”

The teacher, though, ad-

mitted then that this was not within his

ken. “While bans are frequent”,

he advised, “A she cannot be plurized.”

Ne pourrons plus me faire me sen-

tir tout entier. Rester un loup-garou est un sort

qui m’alarme. ” Mais comme le loup était bien poli Il remercia et puis repartit.

The Werewolf knew better–his

sons still slept At home, and homewards now

he crept, Happy, humble, without apolo-

gy For such folly of philology.

The banshee, rising clam-

mily, wailed: “What about my

family?” Then, being not a learned

creature, said humbly “Thanks” and

left the teacher.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Francan de R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

http://www.poezio.net/version?poem-id=825&version-id=1655,1657,1660